OpenStudy (anonymous):

The equation of a circle is (x - 3)2 + (y - 7)2 = 25. Determine the length of the radius. A. 4 B. 25 C. 12.5 D. 5

OpenStudy (anonymous):

d

OpenStudy (anonymous):

Monster!(:

jimthompson5910 (jim_thompson5910):

(x - 3)2 + (y - 7)2 = 25 (x - 3)2 + (y - 7)2 = 5^2 The equation is now in the form (x-h)^2+(y-k)^2=r^2 where r = 5 So the radius is 5 units

OpenStudy (anonymous):

thankyou